OpenStudy (anonymous):

Evaluate the expression described below if the variable given is 9. Two times the difference of one-third of a number and three.

OpenStudy (psymon):

Two times = 2 * _________ Difference of = ____ - and ____ 0ne-third of a number = (1/3)n This is just the clarify what each part of the sentence is telling me. In the first line, I have 2* ______. Well, the blank is the difference of 1/3of a number and 3. So I can start by writing it like this 2(difference of 1/3 number and 3) The second line shows that we have two blanks with a minus in between them. The first part is one-third of a number, which is (1/3)n. Because the phrase says and 3, I replace the and in the second line with 3. So the whole second line I can write as: \[\frac{ 1 }{ 3 }n - 3\] So this result I can multiply by 2 Kind of hard to explain word problems, so definitely ask xD 2(difference of 1/3 number and 3) = \[2(\frac{ 1 }{ 3 }n - 3)\] \[\frac{ 2n }{ 3 }-6\] Now it says the variable is equal to 9, so: \[\frac{ (2)(9) }{ 3 }-6\] \[\frac{ 18 }{ 3 }-6 \] \[6-6= 0\]

OpenStudy (anonymous):

thx